OpenStudy (anonymous):

What could be the value(s) of "x" if the distance between M(2,3) and N(x,15) is 13?

OpenStudy (aakashsudhakar):

So you want to solve for the x-value given several key pieces of information. If D(x, y) = √((x2 - x1)^2 + (y2 - y1)^2) D(x, y) = 13 x1 = 2 y1 = 3 y2 = 15 Find x2. Just use some simple algebra. In this instance, I'll just replace x2 with the variable X to make it easier to notate. 13 = √((X - 2)^2 + (15 - 3)^2) 13 = √((X - 2)^2 + 144) 169 = (X - 2)^2 + 144 25 = (X - 2)^2 (X - 2)^2 = 25 X - 2 = ±5 X = 7 OR X = -3 ...BUT WAIT! Go through it all again logically. What's wrong with one of these values? The problem: distances can't be negative! No one can travel a negative distance (unless you're talking about more intricate physics, but that's a whole different story), therefore the second answer is invalid. This leaves the proper answer to be X = 7.